# Zero-Shot Denoising for Fluorescence Lifetime Imaging Microscopy with Intensity-Guided Learning

### v1

@@ 2 Related Work
In this section, we review prior work on microscopy image denoising methods, covering both traditional and deep learning approaches with a focus on techniques designed for realistic microscopy applications. We also discuss the challenges presented by multi-channel and multimodal microscopy methods in FLIM, an area with very limited existing research. Lastly, we explore zero-shot learning methods, highlighting their potential for microscopy denoising, particularly in cases where limited labeled training data limits the applicability of supervised approaches, and examine how these methodologies could specifically enhance FLIM denoising performance.
@@ 3 Method
Figure 1: Architecture of our zero-shot FLIM denoising framework. (a) The dual-encoder network employs channel-specific encoders to process noisy input data, separately handling intensity-scaled parameters $y_{g}$ and $y_{s}$. Three specialized decoders then generate cleaned outputs: intensity-scaled $y_{g}$, intensity and intensity-scaled $y_{s}$. (b) The pre-trained intensity model utilizes a standard U-Net architecture that was trained on general fluorescence microscopy data as a structural prior that guides the main network’s denoising process.
In this section, we present our novel Zero-Shot Denoising Framework with an Intensity-Guided Learning approach, designed to effectively remove noise while preserving fluorescence lifetime information without additional training on other FLIM components. In contrast to conventional deep learning-based denoising approaches, which typically require paired noisy-clean data for all image channels, the proposed method facilitates denoising in a zero-shot setting by exploiting the inherent correlation between fluorescence intensity images and their corresponding lifetime channels. We first formalize the problem formulation, detail the proposed network architecture, and subsequently describe the training procedure and denoising strategy.
@@ 3.1 Problem Formulation
where $\omega$ represents the angular modulation frequency. The low photon counts and detection in FLIM acquisition introduce photon shot noise and additive thermal noise. However, for simplicity, a noisy FLIM image can be represented as $\mathbf{y}=(y_{g},y_{s})$, where:
@@
$y_{g}$ and $y_{s}$ are the intensity-scaled noisy phase coordinates, $I$ represents the clean fluorescence intensity signal, and $n_{g}$, $n_{s}$ denote the respective noise components.
The goal of FLIM denoising is to estimate $\hat{\mathbf{x}}=(\hat{g}\cdot\hat{I},\hat{s}\cdot\hat{I},\hat{I})$, which closely approximates the clean signal $\mathbf{x}=(g\cdot I,s\cdot I,I)$ while preserving the integrity of the lifetime information encoded in the relationship between $g$ and $s$.
Unlike mixed Poisson-Gaussian noise datasets for confocal and two-photon microscopy, FLIM imaging lacks paired training data comprising noisy acquisitions and their clean reference. To overcome this constraint, we propose a zero-shot framework that learns a noise-adaptive mapping function $f_{\theta}$ operating directly on multi-channel noisy FLIM acquisitions:
$\hat{\mathbf{x}}=f_{\theta}(\mathbf{y},f_{\beta}(y_{I})),$ (3)
where $y_{I}$ is the noisy intensity channel, $f_{\beta}(y_{I})$ represents this intensity channel denoised by a pre-trained intensity denoising model $f_{\beta}$ on mixed Poisson-Gaussian datasets [39], which serves as a structural prior to guide the lifetime components denoising process.
@@ 3.2 Network Architecture
We propose a complete framework consisting of an intensity guidance denoising model and a dual-encoder network for FLIM denoising based on U-Net [29], as illustrated in Fig. 1.
@@ 3.2.1 Pre-trained Intensity Model
The pre-trained intensity model $f_{\beta}$ is based on a three-level U-Net architecture trained using a hybrid loss combining self-supervised Noise2Void loss $\mathcal{L}_{N2V}$ [15] and supervised MSE loss $\mathcal{L}_{MSE}$:
@@
This model provides denoised intensity estimates $f_{\beta}(y_{I})$ that serve as structural priors for the main FLIM denoising network, enabling transfer of knowledge from conventional fluorescence microscopy to FLIM denoising.
@@ 3.2.2 Dual-Encoder FLIM Denoising Network
Our main network processes FLIM channels $y_{g}$ and $y_{s}$ through separate encoder paths. The network utilizes three specialized decoders: two channel-specific decoders that independently reconstruct $\hat{g}\cdot\hat{I}$ and $\hat{s}\cdot\hat{I}$ using features from their respective encoder paths to preserve channel-specific information; and a fusion decoder that produces $\hat{I}$ by leveraging features from both encoder pathways through multi-level feature fusion.
The fusion information and upsampling pathway could be presented as:
@@
where [;][;][ ; ] denotes channel-wise concatenation and $\mathcal{D}^{l}$ represents upsampling operations.
This design ensures that the fusion decoder incorporates comprehensive latent space information, enabling effective intensity-guided reconstruction. By combining features from both $y_{g}$ and $y_{s}$ encoder pathways, the network optimizes information flow from shallow and deep coded features, which enhances the ability of decoders for $g$ and $s$ channels to produce more accurate reconstructions of the FLIM components.
@@ 3.3 Loss Function
Our framework is optimized using a comprehensive joint loss function that addresses multiple aspects of FLIM signal fidelity to restore the images while preserving channel-specific information:
@@
Each component serves a specific purpose. The intensity loss ($\mathcal{L}_{intensity}=\|\hat{I}-f_{\beta}(y_{I})\|^{2}_{2}$) aligns our predicted intensity with the pre-trained denoising output. The fidelity loss ($\mathcal{L}_{fidelity}=\|y_{g}-\hat{g}\cdot\hat{I}\|^{2}_{2}+\|y_{s}-\hat{s}\cdot\hat{I}\|^{2}_{2}$) ensures consistency between inputs and reconstructions to preserve the signal characteristics and limit the optimization ranges for the parameters.
The structure loss ($\mathcal{L}_{structure}=[2-\text{SSIM}(\hat{g}\cdot\hat{I},\hat{I})-\text{SSIM}(\hat{s}\cdot\hat{I},\hat{I})]$) majorly maintains structural coherence. Finally, the total variation loss ($\mathcal{L}_{TV}=[\text{TV}(\hat{g}\cdot\hat{I})+\text{TV}(\hat{s}\cdot\hat{I})]$) reduces noise in relatively uniform areas while preserving sharp edges and preventing over-smoothing of critical details.
@@ 3.4 Zero-Shot Training Strategy
The intensity and the lifetime channels influence each other in the noise distribution but do not follow exactly the same distribution since they encode different biological information during measurement. Accordingly, our zero-shot approach directly processes real-world FLIM data containing natural Poisson-Gaussian mixed photon noise statistics in intensity channels. Rather than learning the distribution of the other channel in a supervised way, it instead draws direct inspiration from the intensity channel.
Our implementation employs an iterative optimization procedure using Adam with an initial learning rate of $1\times 10^{-3}$ and weight decay of $1\times 10^{-5}$. The learning rate is adaptively reduced by a factor of 0.5 when progress plateaus. Each denoising procedure processes randomly cropped $256\times 256$ patches from FLIM images, with optimization continuing for $1k$ to $2k$ iterations to ensure convergence. Compared to conventional deep-learning denoising methods that often require lengthy training phases, our zero-shot approach completes optimization in 10–20 seconds on an NVIDIA GeForce RTX 3090 GPU (24GB VRAM) with CUDA acceleration. This rapid performance significantly benefits time-sensitive biological imaging tasks by minimizing computational overload while maintaining high-quality FLIM reconstructions.
@@ 4 Experiments
Figure 2: Visual comparison of FLIM image denoising methods applied to Ex vivo BPAE cells. Top row: full-field images; bottom row: zoomed-in regions corresponding to the highlighted areas. From left to right: Noisy input, BM3D, DnCNN, AVG 5, and our proposed zero-shot method. The color bar represents fluorescence lifetime values ranging from 1–4 ns. the proposed method in this work demonstrates superior preservation of lifetime information comparable to AVG 5 while achieving effective noise reduction.
In this section, we evaluate the proposed Zero-Shot Denoising Framework for FLIM with Intensity-Guided Learning using empirical data. Our methodology incorporates diverse regions of interest (ROI) across multiple specimens to capture authentic noise variations representative of real-world imaging conditions. Through comprehensive experiments across diverse biological specimens—including In vivo, Ex vivo, and fixed samples with different fluorophores— we assess the efficacy of our approach against state-of-the-art denoising methods on real-world FLIM data.
Given the correlation and distinct distribution patterns across channels of FLIM data and limited dataset availability, we focus our evaluation on the quality of lifetime images and their corresponding $g$ and $s$ components. Furthermore, we analyze the generalization capabilities of our framework in lifetime analysis and perform ablation studies to quantify the contributions of individual loss functions within the methodology.
@@ 4.1.1 Model Pretraining
Prior to application on FLIM data, we pre-trained a U-Net architecture backbone network on diverse fluorescence microscopy images containing mixed Poisson-Gaussian noise [39]. This pretraining strategy enables our model to learn the statistical properties of noise distributions common in fluorescence imaging without requiring paired clean-noisy training data specific to FLIM.
Importantly, our approach utilizes general fluorescence microscopy data rather than FLIM-specific images for pretraining. This methodology allows the network to develop generalizable denoising principles that can be effectively transferred to FLIM data during inference through our zero-shot framework, which helps overcome the limited availability of matched noisy-clean FLIM image pairs for training.
@@ 4.1.2 Data Acquisition
Due to the limited availability of public FLIM data, we evaluated the proposed method on five distinct biological specimens from both published datasets and newly acquired images. These specimens, spanning a range of sample types and fluorescence characteristics, were employed for quantitative evaluation: (1) In vivo Zebrafish: labeled with enhanced green fluorescent protein (EGFP) [40]; (2) Ex vivo Mouse Kidney: tissue sections were stained with DAPI (nuclei), Alexa Fluor 488 wheat germ agglutinin (cell membranes), and Alexa Fluor 568 phalloidin (F-actin) to visualize key structural components [22]; (3) Ex vivo BPAE Cells: bovine pulmonary artery endothelial cells were labeled with DAPI (nuclei), Alexa Fluor 488 phalloidin (F-actin), and MitoTracker Red CMXRos (mitochondria) to enable multicomponent lifetime analysis [38]; (4) Fixed NIH3T3 Cells: fibroblasts were stained with DAPI (nuclei), Alexa Fluor 488 phalloidin (F-actin), and Alexa Fluor 594 biocytin (AFB) to examine subcellular structures; and (5) In vivo Mouse Kidney: intrinsic autofluorescence from kidney tissue was imaged in live animal models to assess performance under physiological conditions [21, 38]. This diverse collection provided a comprehensive basis for quantitative evaluation.
This diverse dataset represents a comprehensive testbed for fluorescence lifetime image denoising algorithms, encompassing both exogenous fluorophores with targeted subcellular localization and endogenous autofluorescence signals across different tissue types. Due to the lack of clean benchmarks for In vivo kidney, we only compare this result to DnCNN [22].
@@ 4.2 Real-World Noise Reduction
The ground truth was established by averaging 5-15 images (AVG 5-15) for each FLIM component on the same regions of interest (ROIs), carefully balancing signal quality with the need to minimize photobleaching and photodamage to biological samples during extended imaging sessions. For each denoising method, we calculated the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) for both $g$ and $s$ components. Additionally, we computed the Absolute Lifetime Error (ALE), which quantifies the percentage deviation of calculated fluorescence lifetime values from the ground truth. Comparative results across all samples are summarized in Tab. 1.
@@
Additionally, qualitative analysis of lifetime images with HSV color mapping (lifetime as Hue and intensity as Value) reveals significant performance differences across methods. As shown in Fig. 2, which displays cellular structures, BM3D partially denoises individual $g$ and $s$ components but significantly distorts lifetime values. Similarly, DnCNN fails to preserve critical information in nuclei regions despite its pre-trained intensity channel denoising, as it cannot account for channel correlation information.
The proposed method, however, effectively preserves morphological details while reducing noise. Fig. 3 provides further evidence of the superiority from this work through in vivo Mouse Kidney imaging. Our approach maintains accurate fluorescence lifetime values across different tissue regions, whereas DnCNN introduces visible bias. Our zero-shot denoising performs consistently across diverse biological samples, highlighting its robustness. By processing all channels simultaneously and leveraging pre-trained intensity maps, our approach effectively preserves essential lifetime information while eliminating noise.

### v2

@@ 2 Related Work
In this section, we review prior work on microscopy image denoising methods, covering both traditional and deep learning approaches with a focus on techniques designed for realistic microscopy applications. We also discuss the challenges presented by multi-channel and multimodal microscopy methods like FLIM, an area with very limited existing research. Lastly, we explore zero-shot learning methods, highlighting their potential for microscopy denoising, particularly in cases where limited labeled training data limits the applicability of supervised approaches, and examine how these methodologies could specifically enhance FLIM denoising performance.
@@ 3 Method
Figure 1: Architecture of our zero-shot FLIM denoising framework. (a) The dual-encoder network employs channel-specific encoders to process noisy input data, separately handling intensity-scaled parameters $y_{g}$ and $y_{s}$. Three specialized decoders then generate cleaned outputs: intensity-scaled $\hat{y_{g}}$, intensity $\hat{y_{I}}$ and intensity-scaled $\hat{y_{s}}$. (b) The pre-trained model utilizes a standard U-Net architecture that was trained on general fluorescence microscopy intensity data $y_{I}$ as a structural prior that guides the main network’s denoising process.
In this section, we introduce our Zero-Shot Denoising Framework with Intensity-Guided Learning. Our framework effectively reduces noise in fluorescence lifetime images across all channels, preserving important fluorescence lifetime details and enhancing image quality. In contrast to conventional deep learning denoising methods that typically rely on paired noisy-clean datasets for each image channel, the proposed method enables denoising in a zero-shot setting by exploiting the inherent correlation between fluorescence intensity images and their corresponding lifetime channels. We first formalize the problem and then describe the proposed network architecture. Afterward, we explain the training procedure and the strategy used for denoising.
@@ 3.1 Problem Formulation
where $\omega$ represents the angular modulation frequency. The low photon counts and detection in FLIM acquisition introduce photon shot noise and additive Gaussian noise. For simplicity, a noisy FLIM image can be represented as $\mathbf{y}=(y_{g},y_{s})$, where:
@@
$y_{g}$ and $y_{s}$ are the intensity-scaled noisy phasor coordinates, $I$ represents the clean fluorescence intensity signal, and $n_{g}$, $n_{s}$ denote the respective noise components.
The goal of FLIM denoising is to estimate the clean signal vector $\hat{\mathbf{y}}=(\hat{y}_{g},\hat{y}_{s},\hat{y}_{I})$, where $\hat{y}_{g}=\hat{I}\cdot\hat{g}$ and $\hat{y}_{s}=\hat{I}\cdot\hat{s}$. Here, $\hat{g}$ and $\hat{s}$ represent the estimated phasor coordinates, and $\hat{I}$ is the estimated fluorescence intensity map. This estimated vector $\hat{\mathbf{y}}$ should closely approximate the clean signal $\mathbf{y^{*}}=(g\cdot I,s\cdot I,I)$, while preserving the accuracy of the lifetime information embedded in the relationship between $\hat{g}$ and $\hat{s}$.
Unlike mixed Poisson-Gaussian noise datasets for confocal and two-photon microscopy, FLIM imaging lacks paired training data comprising noisy images and their clean reference. To overcome this constraint, we make use of a neural network that learns a noise-adaptive mapping function $f_{\theta}$ operating directly on multi-channel noisy FLIM acquisitions in a zero-shot manner to obtain clean vector estimations $\hat{\mathbf{y}}$ :
$\hat{\mathbf{y}}=f_{\theta}(\mathbf{y},f_{\beta}(y_{I})),$ (3)
where $y_{I}$ is the noisy intensity channel, $f_{\beta}(y_{I})$ rrepresents the output of the pre-trained intensity denoising model $f_{\beta}$ trained on mixed Poisson-Gaussian datasets [39], which serves as a structural prior to guide the lifetime components denoising process.
@@ 3.2 Network Architecture
Our complete framework consists of an intensity guidance denoising model and a dual-encoder network for FLIM denoising based on U-Net [29], as illustrated in Fig. 1.
@@ 3.2.1 Pre-trained Intensity Model
The pre-trained intensity model $f_{\beta}$ employs a three-level U-Net architecture optimized using a hybrid loss function. This function combines a self-supervised Noise2Void loss $\mathcal{L}_{N2V}$ [15] with a supervised mean squared error (MSE) loss $\mathcal{L}_{MSE}$:
@@
In the encoder pathway, image features are progressively refined through cascaded double-convolution blocks, with the number of feature channels sequentially increased from $1$ to $32$, $64$, $128$, and finally $256$. Each convolutional layer is followed by a LeakyReLU activation to facilitate robust gradient propagation. Symmetrically, the decoder pathway reconstructs spatial details using transposed convolutions, integrating low-level features from the encoder through dense skip connections. These skip connections ensure fine-grained spatial information is retained. The resulting pre-trained model provides denoised intensity estimates $f_{\beta}(y_{I})$, which act as structural priors for the primary FLIM denoising network, enabling the transfer of knowledge from conventional fluorescence microscopy images to enhance FLIM denoising performance.
@@ 3.2.2 Dual-Encoder FLIM Denoising Network
Our main network processes FLIM channels $y_{g}$ and $y_{s}$ through separate encoder paths, each with two hierarchical levels (1→32→64 channels) using $3\times 3$ kernels, BatchNorm, and ReLU activations. The network utilizes three specialized decoders: two channel-specific decoders that independently reconstruct $\hat{y}_{g}$ and $\hat{y}_{s}$ using features from their respective encoder paths to preserve channel-specific information; and a fusion decoder that produces $\hat{y}_{I}$ by leveraging features from both encoder pathways through multi-level feature fusion.
The network is initialized with He normal weight initialization to facilitate robust zero-shot capability on each sample. The fusion decoder begins by concatenating the deepest features from both encoders and progressively upsamples through transposed convolutions. At each level, skip connections incorporate encoder features from both pathways:
@@
where [;][;][ ; ] denotes channel-wise concatenation and $\mathcal{D}^{l}$ represents upsampling operations followed by double convolution blocks that reduce feature dimensions (128→64→32→1).
This design ensures that the fusion decoder incorporates comprehensive latent space information, enabling effective intensity-guided reconstruction from $y_{g}$ and $y_{s}$. By strategically combining features from both encoder pathways across multiple scales, the network optimizes information flow from both shallow spatial details and deep semantic features. This integration enhances the ability of decoders for the $g$ and $s$ channels to produce more accurate reconstructions of the FLIM components.
@@ 3.3 Loss Function
Our framework is optimized using a comprehensive joint loss function that try to address spatial resolution and cross-channel consistency of FLIM signal fidelity to restore the images. This approach also preserves cross channel information for lifetime accuracy. The loss function combines multiple terms as shown:
@@
The intensity loss ($\mathcal{L}_{intensity}=\|\hat{y}_{I}-f_{\beta}(y_{I})\|^{2}_{2}$) aligns our predicted intensity with the pre-trained denoising output. This serves as the intensity denoising prior for the model optimization. The fidelity loss ($\mathcal{L}_{fidelity}=\|y_{g}-\hat{y}_{g}\|^{2}_{2}+\|y_{s}-\hat{y}_{s}\|^{2}_{2}$) ensures consistency between inputs and reconstructions, thereby preserving the intrinsic signal characteristics. This component effectively constrains the optimization parameters within physiologically plausible bounds, which is particularly crucial as $y_{g}$ and $y_{s}$ encode the lifetime information that should be accurately preserved during the learning process.
The structure loss ($\mathcal{L}_{structure}=[2-\text{SSIM}(\hat{y}_{g},\hat{y}_{I})-\text{SSIM}(\hat{y}_{s},\hat{y}_{I})]$) maintains structural coherence across different reconstruction components. Since the estimated parameters $\hat{y}_{s}$ and $\hat{y}_{g}$ represent intensity scaling $g$ and $s$, their structural information should correlate closely with the estimated intensity $\hat{y}_{I}$ to preserve biologically relevant morphological features in the reconstructed images. Finally, the total variation loss ($\mathcal{L}_{TV}=[\text{TV}(\hat{y}_{g})+\text{TV}(\hat{y}_{s})]$) reduces noise in relatively uniform areas while preserving sharp edges and preventing over-smoothing of critical details. This regularization term is particularly appropriate due to the normalization constraints inherent in the deconvolution process, effectively limiting the parameter space during optimization and mitigating potential overfitting when gradient steps are excessively large.
@@ 3.4 Zero-Shot Training Strategy
The intensity and lifetime channels influence each other in the noise distribution but do not follow exactly the same distribution because they encode different biological information during measurement. Accordingly, our zero-shot approach directly processes real-world FLIM data containing natural Poisson-Gaussian mixed photon noise statistics in the intensity channel. This method is implemented as a test-time optimization per image, which means it functions in a zero-shot manner without requiring any fine-tuning on new samples. Rather than learning the distribution of the other channel in a supervised way, it instead draws direct inspiration from the intensity channel.
Our implementation employs an iterative optimization procedure using Adam with an initial learning rate of $1\times 10^{-3}$ and weight decay of $1\times 10^{-5}$. The learning rate is adaptively reduced by a factor of 0.5 when progress plateaus. Each denoising procedure processes randomly cropped $256\times 256$ patches from FLIM images, with optimization continuing for $1k$ to $2k$ iterations to ensure convergence. Compared to conventional deep-learning denoising methods that often require lengthy training phases, our zero-shot approach completes optimization in 10–20 seconds on an NVIDIA GeForce RTX 3090 GPU (24GB VRAM) with CUDA acceleration per test. This rapid performance significantly benefits time-sensitive biological imaging tasks by minimizing computational overload while maintaining high-quality FLIM reconstructions.
@@ 4 Experiments
Figure 2: Visual comparison of FLIM image denoising methods applied to Ex vivo BPAE cells. Top row: full-field images; bottom row: zoomed-in regions corresponding to the highlighted areas. From left to right: Noisy input, BM3D, DnCNN, Ground Truth(AVG 5), and our proposed zero-shot method. The color bar represents fluorescence lifetime values ranging from 1–4 ns. the proposed method in this work demonstrates superior preservation of lifetime information comparable to AVG 5 while achieving effective noise reduction.
We evaluate the proposed Zero-Shot Denoising Framework for FLIM using empirical data in this section. Our methodology incorporates diverse regions of interest (ROI) across multiple specimens to capture authentic noise variations representative of real-world imaging conditions. Through comprehensive experiments across diverse biological specimens—including In vivo, Ex vivo, and fixed samples with different fluorophores—we assess the efficacy of our approach against state-of-the-art denoising methods on real-world FLIM data.
Given the correlation and distinct distribution patterns across channels of FLIM data, coupled with the inherent limitations in dataset availability, we concentrate our evaluation on two critical aspects: (1) the quality of reconstructed lifetime images and their corresponding $g$ and $s$ components, and (2) the accuracy of lifetime estimations produced by our framework. Our comprehensive analysis demonstrates significant improvements over existing state-of-the-art methods in both qualitative and quantitative metrics. We also conduct comprehensive ablation studies to measure how each loss function uniquely contributes to the overall performance of the model. These experiments reveal important insights about how our different optimization objectives work together.
@@ 4.1.1 Model Pretraining
Prior to application on FLIM data denoising, we apply a pre-trained U-Net architecture backbone network on real-world fluorescence microscopy images containing mixed Poisson-Gaussian noise [39]. This pretraining strategy enables our model to learn the statistical properties of noise distributions common in fluorescence imaging without requiring paired clean-noisy training data specific to FLIM. By utilizing general fluorescence microscopy data for pretraining rather than FLIM-specific images, our network develops generalizable denoising principles that can be effectively transferred to FLIM data during inference through our zero-shot framework. This strategy helps overcome the limited availability of matched noisy-clean FLIM image pairs for training.
@@ 4.1.2 Data Acquisition
Due to the limited availability of public FLIM data, we evaluated the proposed method on five distinct biological specimens from both published datasets and newly acquired images with two-photon excitation at an 800 nm wavelength by “Instant-FLIM” [40]. These specimens, spanning a range of sample types and fluorescence characteristics, were employed for quantitative evaluation: (1) In vivo Zebrafish: labeled with enhanced green fluorescent protein (EGFP) [40]; (2) Ex vivo Mouse Kidney: tissue sections were stained with DAPI (nuclei), Alexa Fluor 488 wheat germ agglutinin (cell membranes), and Alexa Fluor 568 phalloidin (F-actin) to visualize key structural components [22]; (3) Ex vivo BPAE Cells: bovine pulmonary artery endothelial cells were labeled with DAPI (nuclei), Alexa Fluor 488 phalloidin (F-actin), and MitoTracker Red CMXRos (mitochondria) to enable multicomponent lifetime analysis [38]; (4) Fixed NIH3T3 Cells: fibroblasts were stained with DAPI (nuclei), Alexa Fluor 488 phalloidin (F-actin), and Alexa Fluor 594 biocytin (AFB) to examine subcellular structures; and (5) In vivo Mouse Kidney: intrinsic autofluorescence from kidney tissue was imaged in live animal models to assess performance under physiological conditions [21, 38].
Due to the lack of clean benchmarks for (5) In vivo kidney, we only compare this result to DnCNN [22]. This diverse dataset represents a comprehensive testbed for fluorescence lifetime image denoising algorithms, encompassing both exogenous fluorophores with targeted subcellular localization and endogenous autofluorescence signals across different tissue types.
@@ 4.2 Real-World Noise Reduction
The ground truth was established by averaging 5-15 images (AVG 5-15) for each FLIM component on the same ROIs, carefully balancing signal quality with the need to minimize photobleaching and photodamage to biological samples during extended imaging sessions. For each denoising method, we calculated the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) for both $g$ and $s$ components. Additionally, we computed the Absolute Lifetime Error (ALE), which quantifies the percentage deviation of calculated fluorescence lifetime values from the ground truth. Comparative results across all samples with averaged result of $g$ and $s$ are summarized in Tab. 1.
@@
Additionally, qualitative analysis of lifetime images with HSV color mapping (lifetime as Hue and intensity as Value) reveals significant performance differences across methods. As shown in Fig. 2, which displays cellular structures, BM3D partially denoises individual $g$ and $s$ components but significantly distorts lifetime values. Similarly, DnCNN fails to preserve critical information in nuclei regions despite its pre-trained intensity channel denoising, as it cannot account for channel correlation information. The proposed method, however, effectively preserves morphological details while reducing noise.
Fig. 3 provides further evidence of the superiority from this work through in vivo Mouse Kidney imaging. Our approach maintains accurate fluorescence lifetime values across different tissue regions, whereas DnCNN introduces visible bias. Our zero-shot denoising performs consistently across diverse biological samples, highlighting its robustness. By processing all channels simultaneously and leveraging pre-trained intensity maps, our approach effectively preserves essential lifetime information while eliminating noise.